OpenStudy (anonymous):

Simplify the expression: 32-16+2

OpenStudy (anonymous):

Use PEMDAS

OpenStudy (anonymous):

P ( ) E x^2 M x D / A + S -

OpenStudy (anonymous):

How do I simplify my answer 18

OpenStudy (idealist10):

32-16+2 =16+2 =18. Answer: 18

OpenStudy (anonymous):

No So first you will add 16 + 2 Then you will do 32 - 18 Which is 14

OpenStudy (anonymous):

But you have to go left to right

OpenStudy (anonymous):

Yes you are right I'm sorry

OpenStudy (idealist10):

@SCARLET_SPIDER , you're wrong. The answer is 18. Why don't you grab a calculator and calculate it? I used my Ti-89 calculator and it says 18.

OpenStudy (anonymous):

36-16=16 and 16+2 = 18 therefore, the answer is 18

OpenStudy (anonymous):

I said sorry @Idealist10 my bad I just miss calculated

OpenStudy (idealist10):

That's fine. :)